OpenStudy (anonymous):

please help! Write a polynomial function with rational coefficients so that p(x) = 0 has the given roots. 5, 2i

OpenStudy (anonymous):

I really have no idea how to do this. By the way, please explain the best you can! I have a test tomorrow

OpenStudy (anonymous):

anyone there?

OpenStudy (anonymous):

(x-5)(x-2i)(x+2i)= (x-5)(x^2-4i^2)= (x-5)(x^2+4)=x^3+4x-5x^2-20= x^3-5x^2+4x-20 p(x)=x^3-5x^2+4x-20

OpenStudy (dumbcow):

remember complex roots always come in pairs so there are really 3 roots 5,2i, -2i

OpenStudy (anonymous):

why is it 4i^2 in the second part?

OpenStudy (anonymous):

(a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

marina come here please http://openstudy.com/study#/updates/4f434cf9e4b065f388dc9945